We drive a distance of 1 km at 16 km/h. Then we drive an additional distance of 1 km at 32 km/h. What is our average speed?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

speed =distance/time

  t1=d1/v1

  t2=d2/v2

 total time=t1 + t2

 total distance =d1 + d2

 average speed = total distance/total time

                       =64/3km/h